Circular No. 6361

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

PROBABLE BINARY MICROLENSING EVENT The MACHO collaboration (cf. IAUC 6312) reports a likely binary

gravitational microlensing event in progress towards the galactic bulge: “This event was initially reported on Mar. 12 as MACHO Alert 96-BLG-3,

via e-mail and our WWW page (http://darkstar.astro.washington.

edu/). During the last few days, it has shown large deviations from a single-lens microlensing fit, which are well fit by a binary lens. The star in question is located at α = 18^{h}01^{m}45^{s}.5, δ = −29^{o}49′47′′ (equinox 2000.0); it was constant at R = 18.1, V = 19.0 during 1993–1995 and has shown a complex lightcurve since 1996 Mar. 12. The star appears to have undergone a caustic crossing (see Udalski et al. 1994, Ap.J. 436, L103, for an example of this) on Mar. 25, and it is currently at 2.5 mag above normal brightness and rising. The binary-lens fit predicts that a second caustic crossing will occur in the next few days, most probably near Mar. 29.5 UT; this will consist of a sharp drop in brightness by > 1.5 mag within a few hours. Very frequent photometry and spectroscopy (hourly or more often) during the caustic crossing can provide valuable information about the location of the lens and the limb-darkening of the source star. A finding chart and additional information are available by anonymous

FTP at darkstar.astro.washington.edu:/macho/Alerts/96-BLG-3 or

at the above WWW site. Potential observers are requested to contact macho@astro.washington.edu, D. Bennett at telephone 510-423-0656, or S. Marshall at 510-422-4872, to coordinate observations.”

COMET C/1995 O1 (HALE-BOPP) A. Fitzsimmons and I. M. Cartwright, Queen’s University of Belfast;

and D. Pollacco, Royal Greenwich Observatory, communicate: “An optical spectrum (range 360–520 nm) of comet C/1995 O1 was obtained using the 2.5-m Isaac Newton Telescope on La Palma on Mar. 19.264 UT. CN(0-0) 388-nm emission was observed extending to 45′′ from the nucleus, equivalent to 1.7 × 10^{5} km at the comet. Assuming HCN as a parent, the preliminary CN production rate is 1.1 × 10^{26}, indicating an increase by a factor of 2 from earlier observations (cf. IAUC 6252), and similar to the increase in the CO-production rate reported on IAUC 6312. No other emission features were observed.”

1996 March 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green